# Mergers Fall Short: Non-merger Channels Required for Galactic Heavy Element Production

### v2

@@ Abstract
Since the discovery of the binary neutron star merger GW170817 and its associated kilonova, neutron star mergers have been established as a key production channel for -process elements in the Universe. However, various evidences, including the observations of -process abundances as inferred from stellar spectra of Milky Way disk stars, suggest that additional channels are needed to fully account for the -process element enrichment in the Milky Way. Neutron star–black hole mergers and fast-merging binary neutron star systems are among the leading alternative candidates. In this paper, we combine gravitational-wave observations from LIGO–Virgo–KAGRA with data from short gamma-ray bursts, Galactic pulsars, and Galactic [Eu/Fe] versus [Fe/H] abundance observations to assess the contribution of these mergers to -process enrichment in the Galactic disk. We find that neither neutron star–black hole mergers nor fast-merging binary neutron star populations can serve as the dominant additional channel without generating strong tension with existing observations and theoretical expectations. These results constrain the viable sources of Galactic -process enrichment and underscore the necessity of non-merger production channels.
@@ I Introduction
The origins of the heaviest elements in the Universe via rapid neutron capture (-process) nucleosynthesis remains a key question in astrophysics. While it is known that the -process happens in neutron-rich environments Burbidge et al. (1957); Cameron (1957), their dominant astrophysical sites are still debated. Compact binary mergers, particularly binary neutron star (BNS) mergers, have been recognized as promising sources of -process elements Lattimer and Schramm (1974); Eichler et al. (1989); Freiburghaus et al. (1999). The detection of GW170817 and its associated kilonova provided direct evidence for -process nucleosynthesis in BNS mergers Abbott and Collaboration) (2017); Abbott et al. (2017); Kasen et al. (2017). However, the relative role of BNS mergers compared to other potential channels, such as neutron star-black hole (NSBH) mergers Rosswog (2005); Kyutoku et al. (2015), magnetorotational supernovae Winteler (2012); Nishimura et al. (2015), collapsars Siegel et al. (2019a), and giant flares from magnetars Cehula et al. (2024); Patel et al. (2025) is an open problem. For example, recent chemical evolution studies have raised concerns about whether the event rate evolution of BNS mergers are consistent with the early enrichment seen in metal-poor stars in the Milky Way Côté (2019); Hotokezaka et al. (2018); van de Voort et al. (2015); Cescutti et al. (2015); van de Voort et al. (2022); Kobayashi et al. (2023); Kolborg et al. (2023). Similar concerns and conclusions are derived from chemical evolution studies of (ultra-faint) dwarf galaxies and globular clusters Bonetti et al. (2019); Zevin et al. (2019); Skúladóttir et al. (2019); Naidu et al. (2022); Simon et al. (2023); Kirby et al. (2023).
@@
One of the observational probes of -process enrichment in the Milky Way is the relationship between and (where ; denotes number abundances and denotes solar), which traces the evolution of europium (Eu, a representative -process element) relative to iron (Fe) as the Galaxy chemically evolves. Observations of metal-poor stars, particularly in the Galactic halo and thick disk, reveal a significant presence of Eu even at low Sneden et al. (2008); Frebel (2010), indicating that -process production was already occurring in the early history of the Galaxy. In order to match the early enrichment, at least some events must occur promptly after star formation—an argument supported by stars like J15213538, an extremely -process-enhanced star with and Cain et al. (2020). The relative abundance of Eu decreases at higher metallicities, implying an increase in late-time Fe production and/or decrease in Eu enrichment. These trends—particularly the negative slope in the [Eu/Fe] versus [Fe/H] relation observed in Milky Way disk stars—place stringent constraints on the rates and yields (mass of -process elements produced per event) of -process sources throughout cosmic history. Using – evolution as a reference, one can test whether the proposed astrophysical sites can reproduce the chemical evolution history of the Milky Way.
In our recent study Chen et al. (2025), we investigated whether BNS mergers alone can account for the observed -process enrichment of disk stars in the Milky Way ([Fe/H]). Standard BNS (”delayed BNS” henceforth) formation scenario predicts a delay relative to the star formation rate (SFR) Peters (1964); Beniamini and Piran (2019). We modeled Eu production from BNS mergers using a one-zone galactic chemical evolution model Siegel et al. (2019b) that incorporates models of star formation history from X-ray observations Madau and Fragos (2017), delay-time distributions constrained by short gamma-ray burst (sGRB) observations Zevin et al. (2022), event rates and mass distribution constrained by gravitational-wave (GW) observations Abbott et al. (2023a, b), neutron star equation-of-state (EOS) constraints by GW and pulsar observations Legred et al. (2021), and semi-analytical fits to numerical simulations of BNS mergers for event yields Krüger and Foucart (2020). The study found that, to explain the observed abundance trends, BNS mergers would need to occur with a delay-time distribution of power-law index and minimum delay time Myr at 90% confidence (following the delay-time model in Eq. 2 of Zevin et al. (2022) ), contrasting those inferred from sGRBs Wanderman and Piran (2015); Ghirlanda et al. (2016); Zevin et al. (2022). If BNS mergers were assumed to follow the sGRB-inferred delays, an additional enrichment channel is necessary. The study concluded that this additional channel must track the SFR closely and likely contribute a dominant part ( at 90% confidence) of the Eu yield in the Galaxy. Similar conclusions were also found in Côté (2019); Kobayashi et al. (2023).
Following these conclusions, it becomes essential to explore other potential contributors to the -process. NSBH mergers and fast-merging BNS are two strong candidates among the discussion. There have been studies on -process contributions from NSBH systems under various mass and spin configurations Foucart et al. (2018); Mapelli et al. (2018). Similarly, fast-merging BNS, potentially originating from dynamical interactions in dense stellar environments such as globular clusters or galactic nuclei Samsing et al. (2018); Rodriguez et al. (2018a), may exhibit distinct ejecta properties and shorter merger timescales compared to field BNSs Vigna-Gómez et al. (2018). Both channels have been speculated as potential candidates to explain the shortfall in the observed abundance evolution, particularly under the assumption that they are either prompt and/or more efficient ejectors Safarzadeh et al. (2019); Bonetti et al. (2019); Andrews and Zezas (2019); Beniamini and Piran (2024); Nugent et al. (2024); Holmbeck and Andrews (2024); Maoz and Nakar (2025). Our goal here is to do a rigorous data-supported assessment of the viability of these two channels to fill in the missing components in the origin of heavy elements in the Milky Way.
Using a likelihood-based inference framework proposed in Chen et al. (2025), we quantify the extent to which the NSBH and fast-merging BNS channels can contribute to the observed evolution of [Eu/Fe] versus [Fe/H]. In doing so, we also examine whether the required conditions are astrophysically plausible given current constraints on merger rates, delay times, and ejecta properties.
@@ II.1 Chemical Evolution Framework with Multiple -process Channels
Our chemical evolution model, incorporating multiple candidate -process enrichment channels, builds directly upon the one-zone prescription employed in Chen et al. (2025), which in turn follows the approach of Siegel et al. (2019b). In this framework, the Milky Way is treated as a single, well-mixed reservoir of interstellar medium (ISM). The model evolves the abundances of heavy elements over cosmic time by integrating nucleosynthesis yields from astrophysical events convolved with their respective rate over time. If the astrophysical events were delayed relative to the SFR, we adopt a delay-time model, which is a power law distribution with minimum delay time () and slope () as parameters (e.g. Eq. 2 in Zevin et al. (2022)). For a given SFR, delay-time model, and event yields, the model tracks the enrichment history in both and iron-peak elements, as well as -process elements such as Eu.
Key enrichment channels in the baseline model include core-collapse supernovae (CCSNe) and Type Ia supernovae (SNe Ia), mainly contributing to and iron enrichment, respectively. The -process enrichment is primarily attributed to delayed BNS mergers, but the model is flexible enough to include additional candidate channels such as fast-merging BNS systems and NSBH mergers, as we explore in this work. Each -process channel is described by its local event rate, the average event yield, a delay-time model or one that directly follows the SFR without additional parameters needed.
@@ II.2 Likelihood Model and Bayesian Inference Framework
The chemical evolution model described in previous subsection generates a time-ordered track in the same abundance space, and the likelihood for each star is computed by integrating its Gaussian distribution over the model track. The total likelihood is the product over all stars, evaluated numerically. We use this likelihood for Bayesian inference on the parameters governing each -process channel, including the local event rate, yield, and, where applicable, the minimum delay time () and delay-time distribution power-law index (). Priors on the local merger rates as a function of binary mass and event yield are formulated from GW-inferred population properties of GWTC-4.0 catalog The LIGO Scientific Collaboration, the Virgo Collaboration, the KAGRA Collaboration (2025), pulsar observations Legred et al. (2021), and semi-analytical fits to numerical simulations Krüger and Foucart (2020), while priors on the delay-time distribution parameters are obtained from sGRB constraints Zevin et al. (2022). We use the Dynesty sampler via the Bilby inference framework to sample from the likelihood surface Speagle (2020); Ashton et al. (2019).
@@ III.1 Current GW searches can detect fast-merging BNSs
We assume that all BNS populations share the same neutron star EOS and model of event yield Krüger and Foucart (2020); Chen et al. (2025).
Our inference shows that reproducing the observed [Eu/Fe] trend in this scenario requires an exceptionally high fast-merging fraction. The posterior for in the local Universe () peaks near unity, with more than of BNS mergers inferred to be fast-merging (90% confidence lower limit). This fraction goes up to at and at , as a SFR-tracking channel is expected to contribute more frequently in the early Universe compared to delayed sources. In Figure 1, we show the corresponding volumetric rate as a function of redshift. In particular, the local () volumetric rate of the fast-merging BNS is Gpc while the delayed BNS rate remains at a lower level of Gpc, where the error bars correspond to 90% credible levels (The reader may refer to Appendix and Fig. 4 for the posteriors of all inferred parameters.). Since this analysis assumes that GW searches are sensitive to both channels, the sum of the two channels is limited by the GW-inferred local BNS rate (). It will not be the case for our analysis in the next subsection.
@@ III.2 Current GW searches miss fast-merging BNSs
Our analysis finds that, in order to explain the observed [Eu/Fe] versus [Fe/H] trend (negative slope for disk stars, as mentioned before) with only delayed and undetected fast-merging BNSs, the multiplicative factor is estimated to be in the local Universe (90% confidence lower limit), while it further rises up to at and at . Equivalently, the fast-merging BNS rate would be Gpc in the local Universe (Figure 2, 90% confidence lower limit), and goes up to Gpc at and Gpc at . This implies that the Milky Way would require roughly 9 times (median value) as many fast-merging BNSs as delayed ones; this outcome again challenges the current understanding of the delayed and fast-merging BNS rates (Beniamini and Piran, 2019; Belczynski et al., 2018; Giacobbo and Mapelli, 2018) (The reader may refer to Appendix and Fig. 5 for the posteriors of all inferred parameters.).
@@ III.3 NSBH Mergers as a Secondary Channel
Figure 3 shows the inferred (rate yield) fraction integrated over cosmic history until present day, from the BNS (blue), NSBH (orange), and third (green) channels. The BNS contribution is consistent with the findings of Chen et al. (2025), while the NSBH contribution remains marginal. In contrast, the third, phenomenological SFR-tracking channel dominates the posterior. Quantitatively, at 90% confidence, more than 70% of the contribution must be from the unknown third channel while the combined contribution from BNS and NSBH is limited to below 30% (The reader may refer to Appendix and Fig.. 6 for the posteriors of all inferred parameters.). These results underscore that BNS and NSBH mergers alone cannot account for the observed [Eu/Fe] versus [Fe/H] trend, and that a substantial additional source of -process elements is required.

### v3

@@ Abstract
Since the discovery of the binary neutron star merger GW170817 and its associated kilonova, neutron star mergers have been established as a key production channel for -process elements in the Universe. However, various evidences, including the observations of -process abundances as inferred from stellar spectra of Milky Way disk stars, suggest that additional channels are needed to fully account for the -process element enrichment in the Milky Way. Neutron star–black hole mergers and fast-merging binary neutron star systems are among the leading alternative candidates. In this paper, we combine gravitational-wave observations from LIGO–Virgo–KAGRA with data from short gamma-ray bursts, Galactic pulsars, and Galactic [Eu/Fe] versus [Fe/H] abundance observations to assess the contribution of these mergers to -process enrichment in the Galactic disk. Our analysis employs a unified, likelihood-based inference framework that consistently propagates uncertainties in merger rates, delay-time distributions, mass- and spin-dependent ejecta yields, and stellar abundance measurements. We find that neither neutron star–black hole mergers nor fast-merging binary neutron star populations can serve as the dominant additional channel without generating strong tension with existing observations and theoretical expectations. These results constrain the viable sources of Galactic -process enrichment and underscore the necessity of non-merger production channels.
@@ I Introduction
The origins of the heaviest elements in the Universe via rapid neutron capture (-process) nucleosynthesis remains a key question in astrophysics. While it is known that the -process happens in neutron-rich environments Burbidge et al. (1957); Cameron (1957), their dominant astrophysical sites are still debated Cowan et al. (2021); Siegel (2022). Compact binary mergers, particularly binary neutron star (BNS) mergers, have been recognized as promising sources of -process elements Lattimer and Schramm (1974); Eichler et al. (1989); Freiburghaus et al. (1999). The detection of GW170817 and its associated kilonova provided direct evidence for -process nucleosynthesis in BNS mergers Abbott and Collaboration) (2017); Abbott et al. (2017); Kasen et al. (2017). However, the relative role of BNS mergers compared to other potential channels, such as neutron star-black hole (NSBH) mergers Rosswog (2005); Kyutoku et al. (2015), magnetorotational supernovae Winteler (2012); Nishimura et al. (2015), collapsars Siegel et al. (2019a), and giant flares from magnetars Cehula et al. (2024); Patel et al. (2025) is an open problem. For example, recent chemical evolution studies have raised concerns about whether the event rate evolution of BNS mergers are consistent with the early enrichment seen in metal-poor stars in the Milky Way Côté (2019); Hotokezaka et al. (2018); van de Voort et al. (2015); Cescutti et al. (2015); Kobayashi et al. (2020); van de Voort et al. (2022); Kobayashi et al. (2023); Kolborg et al. (2023). Similar concerns and conclusions are derived from chemical evolution studies of (ultra-faint) dwarf galaxies and globular clusters Bonetti et al. (2019); Zevin et al. (2019); Skúladóttir et al. (2019); Naidu et al. (2022); Simon et al. (2023); Kirby et al. (2023).
@@
A key observational probe of -process enrichment in the Milky Way is the relationship between and (where ; denotes number abundances and denotes solar), which traces the evolution of europium (Eu, a representative -process element) relative to iron (Fe) as the Galaxy chemically evolves. Observations of metal-poor stars, particularly in the Galactic halo and thick disk, reveal a significant presence of Eu even at low Sneden et al. (2008); Frebel (2010), indicating that -process production was already occurring in the early history of the Galaxy. In order to match the early enrichment, at least some events must occur promptly after star formation—an argument supported by stars like J15213538, an extremely -process-enhanced star with and Cain et al. (2020). The relative abundance of Eu decreases at higher metallicities, implying an increase in late-time Fe production and/or decrease in Eu enrichment. These trends—particularly the negative slope in the [Eu/Fe] versus [Fe/H] relation observed in Milky Way disk stars—place stringent constraints on the rates and yields of -process sources throughout cosmic history. Here, by “yield” we mean the per-event -process yield, i.e. the total mass of -process material that a single event ejects into the Galactic interstellar medium (often referred to as the nucleosynthesis or chemical yield). Using – evolution as a reference, one can test whether the proposed astrophysical sites can reproduce the chemical evolution history of the Milky Way.
In our recent study Chen et al. (2025), we investigated whether BNS mergers alone can account for the observed -process enrichment of disk stars in the Milky Way ([Fe/H]). Standard BNS (“delayed BNS”) formation scenarios predict a delay relative to the star formation rate (SFR) Peters (1964); Beniamini and Piran (2019). Using a one-zone chemical evolution model Siegel et al. (2019b) informed by cosmic SFR models Madau and Fragos (2017), short GRB (sGRB)-inferred delay-time distributions Zevin et al. (2022), GW-inferred BNS rates and mass distributions Abbott et al. (2023a, b), neutron star EOS constraints Legred et al. (2021), and semi-analytical ejecta fits Krüger and Foucart (2020), we found that BNS mergers following the sGRB-inferred delays cannot reproduce the observed disk-star [Eu/Fe] trend. An additional enrichment channel that tracks the SFR is therefore required, consistent with similar conclusions in the literature Côté (2019); Kobayashi et al. (2020, 2023). We note that some studies on the metal-poor halo stars and dwarf galaxies have emphasized the importance of delayed r-process sources in ultra-low-metallicity environments (e.g., Duggan et al. (2018); Tarumi et al. (2021)). These works probe the earliest phases of chemical enrichment, whereas our analysis focuses on the [Eu/Fe] evolution of Milky Way disk stars at higher metallicities.
Given the strong evidence for a prompt enrichment channel, we investigate whether two commonly proposed merger-based candidates—fast-merging BNS systems and NSBH mergers—can plausibly supply the needed early -process production. Fast-merging BNS may originate from dynamical interactions in dense stellar environments such as globular clusters or galactic nuclei Samsing et al. (2018); Rodriguez et al. (2018a), and are often hypothesized to merge promptly or eject more -process material Vigna-Gómez et al. (2018). NSBH mergers have similarly been proposed as efficient ejectors under certain mass and spin configurations Foucart et al. (2018); Mapelli et al. (2018). Both channels have been suggested as possible solutions to the early-enrichment problem, yet it remains unclear whether either can in fact produce the rates and yields required by the observed [Eu/Fe]–[Fe/H] trends. This is the primary question we resolve here.
Using a likelihood-based inference framework proposed in Chen et al. (2025), we assess whether NSBH and fast-merging BNS channels can contribute sufficiently to match the observed evolution of [Eu/Fe] versus [Fe/H], and whether the required conditions are astrophysically plausible given current observational constraints. While the point-estimate studies have shown that standard BNS mergers are unlikely to dominate early Galactic r-process enrichment, such estimates rely on representative values for merger rates, delay times, and ejecta yields. In reality, the r-process yield of compact-binary mergers depends sensitively on the component masses, spins, and equation of state, and the underlying merger population is only weakly constrained by current observations. A likelihood-based inference framework is therefore required to propagate these uncertainties consistently and to determine whether any plausible region of this multi-dimensional parameter space allows merger-based channels to dominate early enrichment. Moreover, point estimates cannot quantify how extreme the required conditions are—for example, whether an additional prompt channel must contribute a modest fraction or more than 95% of the total r-process budget. By marginalizing over population, yield, and observational uncertainties, our approach provides a quantitative assessment of whether fast-merging BNS or NSBH systems can plausibly serve as the missing prompt enrichment source.
@@ II.1 Chemical Evolution Framework with Multiple -process Channels
Our chemical evolution model, incorporating multiple candidate -process enrichment channels, builds directly upon the one-zone prescription employed in Chen et al. (2025), which in turn follows the approach of Siegel et al. (2019b). In this framework, the Milky Way is treated as a single, well-mixed reservoir of interstellar medium (ISM). The model evolves the abundances of heavy elements over cosmic time by integrating the -process yields from astrophysical events convolved with their respective rate over time. If the astrophysical events were delayed relative to the SFR, we adopt a delay-time model, which is a power law distribution with minimum delay time () and slope () as parameters (e.g. Eq. 2 in Zevin et al. (2022)). For a given SFR, delay-time model, and yields, the model tracks the enrichment history in both and iron-peak elements, as well as -process elements such as Eu.
Key enrichment channels in the baseline model include core-collapse supernovae (CCSNe) and Type Ia supernovae (SNe Ia), mainly contributing to and iron enrichment. The -process enrichment is primarily attributed to delayed BNS mergers, but the model is flexible enough to include additional candidate channels such as fast-merging BNS systems and NSBH mergers, as we explore in this work. Each channel is parameterized by a local event rate, per-event -process yield or simply yield, and an evolutionary prescription: it either tracks the SFR directly or is obtained by convolving the SFR with a channel-specific delay–time distribution. For comparison with the observed [Eu/Fe]–[Fe/H] trends, we convert the total -process yield of each event into a Eu yield by assuming a fixed -process abundance pattern and adopting the solar -process Eu mass fraction. Thus, unless explicitly stated otherwise, “yield” refers to the total -process mass, and Eu yields are obtained as a fixed fraction of this quantity. In all cases, the underlying SFR is taken to be the cosmic star-formation history of Madau and Fragos (2017).
@@ II.2 Likelihood Model and Bayesian Inference Framework
The chemical evolution model described in previous subsection generates a time-ordered track in the same abundance space, and the likelihood for each star is computed by integrating its Gaussian distribution over the model track. The total likelihood is the product over all stars, evaluated numerically. We use this likelihood for Bayesian inference on the parameters governing each -process channel, including the local event rate, yield, and, where applicable, the minimum delay time () and delay-time distribution power-law index ().
For the local merger rates and r-process yields of BNS and NSBH systems, we begin with population inference results of component masses and local merger rates as obtained from the analysis of GWTC-4.0 The LIGO Scientific Collaboration, the Virgo Collaboration, the KAGRA Collaboration (2025). These population samples are propagated through the framework of Chen et al. (2021) to obtain joint constraints for the local merger rates and the ejecta masses. In this procedure, neutron-star radii and compactnesses are derived by combining the component masses with independent EOS constraints from joint GW and pulsar observations Legred et al. (2021). The resulting compactnesses are then used as inputs to NR-calibrated semi-analytical fitting formulae Krüger and Foucart (2020) to compute the dynamical ejecta and remnant disk masses with uncertainties as specified in Eq. (1) of Ref. Chen et al. (2025), which are subsequently combined to estimate the total ejecta mass. The resulting joint posterior distributions for the local merger rate and ejecta mass are considered as informative priors in our inference framework.
@@ III.1 Current GW searches can detect fast-merging BNSs
We assume that all BNS populations share the same neutron star EOS and follow the same prescriptions to calculate the -process yields given the component masses Krüger and Foucart (2020); Chen et al. (2025).
Our inference shows that reproducing the observed [Eu/Fe] trend in this scenario requires an exceptionally high fast-merging fraction. The posterior for in the local Universe () peaks near unity, with more than of BNS mergers inferred to be fast-merging (90% confidence lower limit). This fraction goes further up at higher redshifts, as a SFR-tracking channel is expected to contribute more frequently in the early Universe compared to delayed sources. In Figure 1, we show the corresponding volumetric rate as a function of redshift. In particular, the local () volumetric rate of the fast-merging BNS is Gpc while the delayed BNS rate remains at a lower level of Gpc, where the error bars correspond to 90% credible levels (The reader may refer to the Appendix and Fig. 4 for the posteriors of all inferred parameters.). Since this analysis assumes that GW searches are sensitive to both channels, the sum of the two channels is limited by the GW-inferred local BNS rate (). It is not the case for our analysis in the next subsection.
@@ III.2 Current GW searches miss fast-merging BNSs
Our analysis finds that, in order to explain the observed [Eu/Fe] versus [Fe/H] trend (negative slope for disk stars, as mentioned before) with only delayed and undetected fast-merging BNSs, the multiplicative factor is estimated to be in the local Universe (90% confidence lower limit), while it further rises up to at and at . Equivalently, the fast-merging BNS rate would be Gpc in the local Universe (Figure 2, 90% confidence lower limit), and goes up to Gpc at and Gpc at . This implies that the Milky Way would require roughly 17 times (median value) as many fast-merging BNSs as delayed ones at ; this outcome again challenges the current understanding of the delayed and fast-merging BNS rates (Beniamini and Piran, 2019; Belczynski et al., 2018; Giacobbo and Mapelli, 2018) (see the Appendix and Fig. 5 for the posteriors of all inferred parameters.).
@@
The posterior distributions for and , and the corresponding fast-merging BNS rates derived above quantify the event rates that would be required for fast-merging systems to reproduce the observed Galactic [Eu/Fe] trends within our chemical evolution model. They do not, by themselves, represent the probability that such fast-merging populations exist in nature. A fully self-contained Bayesian hypothesis test of fast-merging–dominated enrichment would require well-calibrated external priors on the intrinsic fast-merging BNS rate, for example from population synthesis, short-GRB statistics, or detailed modeling of dense stellar environments. At present, we lack any widely-accepted quantitative constraints on the fast-merging subpopulation in a form that can be robustly incorporated into a second-stage Bayesian model comparison. Our assessment therefore rests on the tension between the exceptionally high fast-merging rates inferred here and current theoretical and observational expectations for compact-binary populations, rather than on a formal statistical exclusion based on externally imposed rate priors.
@@ III.3 NSBH Mergers as a Secondary Channel
Figure 3 shows the inferred (rate yield) fraction integrated over cosmic history until present day, from the BNS (blue), NSBH (orange), and third (green) channels. The BNS contribution is consistent with the findings of Chen et al. (2025), while the NSBH contribution remains marginal. In contrast, the third, phenomenological SFR-tracking channel dominates the posterior. Quantitatively, at 90% confidence, more than 90% of the contribution must be from the unknown third channel while the combined contribution from BNS and NSBH is limited to below 10% (The reader may refer to the Appendix and Fig.. 6 for the posteriors of all inferred parameters.). These results underscore that BNS and NSBH mergers alone cannot account for the observed [Eu/Fe] versus [Fe/H] trend, and that a substantial additional source of -process elements is required.